\begin{thebibliography}{20}
\bibitem{TSP} AA.VV. - \emph{The Stacks Project}. \url{https://stacks.math.columbia.edu/}.
\bibitem{B:B:G} I. Biswas, U. Bruzzo, S. Gurjar - \emph{Higgs bundles and fundamental group schemes}, Adv. Geom. \textbf{19} (2019) 381--388.
\bibitem{B:E} E. Bombieri - \emph{Canonical models of surfaces of general type}, Inst. Hautes \'{E}tudes Sci. Publ. Math. \textbf{42} (1973) 171--219.
\bibitem{B:C:GO} U. Bruzzo, A. Capasso, B. Gra\~na Otero - \emph{Positivity for Higgs vector bundles: criteria and applications}, \url{arXiv:2307.16578v2} \textcolor{blue}{\texttt{[math.AG]}}.
\bibitem{B:GO:1} U. Bruzzo, B. Gra\~na Otero - \emph{Numerically flat Higgs vector bundles}, Commun. Contemp. Math. \textbf{9} (2007) 437--446.
\bibitem{B:GO:2} U. Bruzzo, B. Gra\~na Otero - \emph{Semistable and numerically effective principal (Higgs) bundles}, Advances in Mathematics \textbf{226} (2011) 3655--3676.
\bibitem{B:HR} U. Bruzzo, D. Hern\'andez Ruip\'erez - \emph{Semistability vs. nefness for (Higgs) vector bundles}, Diff. Geom. Appl. \textbf{24} (2006) 403--416.
%\bibitem{B:LG} U. Bruzzo, A. Lo Giudice - \emph{Restricting Higgs bundles to curves}, Asian J. Math. \textbf{20} (2016) 399--408.
%\bibitem{C:P:T} F. Campana, T. Peternell, M. Tama - \emph{Geometric stability of the cotangent bundle and the universal cover of a projective manifold}, Bull. Soc. math. France \textbf{139} (2011) 41--74.
\bibitem{A:C} A. Capasso - \emph{Lefschetz principle-type theorems for curve semistable Higgs sheaves and applications} \url{arXiv:2412.00439v2} \textcolor{blue}{\texttt{[math.AG]}}.
\bibitem{D:O:1} O. Debarre - \emph{Varieties with Ample Cotangent Bundle}, \url{arxiv:0306066v5} \textcolor{blue}{\texttt{[math.AG]}}.
\bibitem{D:O:2} O. Debarre (2001) \emph{Higher-Dimensional Algebraic Geometry}, Springer.
\bibitem{DiC:LF} L. F. Di Cerbo - \emph{The Toledo invariant, and Seshadri constants of fake projective planes}, J. Math. Soc. Japan \textbf{69} (2017) 1601-1610.
\bibitem{F:W} W. Fulton (1998) \emph{Intersection theory. Second edition}, Springer.
\bibitem{GKPT} D. Greb, S. Kebekus, T. Peternell, B. Taji - \emph{The Miyaoka-Yau inequality and uniformisation of canonical
models}, Ann. Scient. Ec. Norm. Sup. \textbf{52} (2019) 1487--1535.
%\bibitem{G:O} F. Gounelas, J. C. Ottem - \emph{Remarks on the positivity of the cotangent bundle of a K3 surface}, Epj.G.A. \textbf{4} (2020) article ID 8.
\bibitem{H:RC} R. C. Hartshorne (1977) \emph{Algebraic Geometry}, Springer.
\bibitem{K:JH} J.H. Keum - \emph{Quotients of fake projective planes}, Geometry \& Topology \textbf{12} (2008) 2497--2515.
\bibitem{L:LG} V. Lanza, A. Lo Giudice, \emph{Bruzzo's conjecture}, J. Geom. Phys. \textbf{118} (2017) 181--191.
\bibitem{L:A} A. Langer - \emph{Bogomolov's inequality for Higgs sheaves in positive characteristic}, Inv. Math. \textbf{199} (2015) 889--920.
\bibitem{L:RK} R. K. Lazarsfeld (2004) \emph{Positivity in algebraic geometry. I and II}, Springer.
\bibitem{M:R} V. B. Mehta, A. Ramanathan - \emph{Semistable sheaves on projective varieties and their restrictions to curves}, Math. Ann. \textbf{258} (1982) 213--224.
\bibitem{M:Y:2} Y. Miyaoka - \emph{On the Chern numbers of surfaces of general type}, Inventiones Mathematicae \textbf{42} (1977) 225--237.
\bibitem{M:Y:3} Y. Miyaoka - \emph{Algebraic surfaces with positive indices}, in K. Ueno ed., Classification of algebraic and analytic manifolds, Katata (1982) 281--301. Progr. Math. \textbf{39}, Birk\"auser (1983) Boston MA.
\bibitem{M:Y:1} Y. Miyaoka - \emph{The Chern classes and Kodaira dimension of a minimal variety}, Algebraic geometry, Sendai (1985) 449--476. Adv. Stud. Pure Math. \textbf{10}, North-Holland Publishing Co. (1987) Amsterdam.
\bibitem{S:CT} C. T. Simpson - \emph{Constructing variations of Hodge structure using Yang-Mills theory and applications to uniformization}, J. Am. Math. Soc. \textbf{1} (1988) 867--918.
\bibitem{T:H} H. Tsuji - \emph{Stability of the tangent bundles of minimal algebraic varieties}, Topology \textbf{27} (1988) 429--442.
\bibitem{Y:S-T} S.-T. Yau - \emph{Calabi's conjecture and some new results in algebraic geometry}, Proc. Nat. Acad. Sci. \textbf{74} (1977) 1798--1799.
\bibitem{FOAG} R. D. Vakil - \emph{Foundations of Algebraic Geometry}. Stanford University, February 21st 2024 \url{https://math.stanford.edu/~vakil/216blog/FOAGfeb2124public.pdf}.
\bibitem{Z:L} M. G. Za\v\i denberg, V. Ya. Lin - \emph{Finiteness theorems for holomorphic mappings}. In G. M. Khenkin ed., Several Complex Variables III, Encyclopedia of Math. Sciences, vol. 9 (1989) 113--172. Springer, Berlin, Heidelberg.
\end{thebibliography}

\documentclass[a4,twoside,10pt]{amsart}

\usepackage{a4wide}
\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage[english]{babel}
\usepackage{enumerate}
\usepackage{fontenc}
\usepackage[colorlinks=true,linkcolor=blue,citecolor=red,urlcolor=blue]{hyperref}
\usepackage[latin1]{inputenc}
\usepackage{orcidlink}
\usepackage{stackrel}
\usepackage[english]{varioref}
\usepackage[all]{xy}
\usepackage{url}

\title{Curve semistable Higgs bundles and smooth projective varieties whose canonical bundle is ample}

\author{Armando Capasso}
\address{Universit\`a degli Studi di Trieste, P.le Europa 1, Trieste (Italy), C.A.P. 34127}
\email{armando.capasso@units.it}

\thanks{A.C. is member of INdAM-GNSAGA. ORCID: 0009-0001-5463-7221 \orcidlink{0009-0001-5463-7221}}

\subjclass[2020]{14D07, 14F05, 14J29, 14J60, 32Q45}
\keywords{Higgs bundles, semistability, curve semistablity, ampleness, numerical effectiveness, variety of general type, Gugghenheimer-Yau inequality, algebraic hyperbolic}

\date{\today}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{propositions}[theorem]{Propositions}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{ex}[theorem]{Example}
\newenvironment{example}{\begin{ex}}{\hfill{$\triangle$}\end{ex}}
\newtheorem{rem}[theorem]{Remark}
\newenvironment{remark}{\begin{rem}}{\hfill{$\Diamond$}\end{rem}}
\newtheorem{remarks}[theorem]{Remarks}
\newtheorem*{Proof}{Proof}
\newenvironment{prf}{\begin{Proof}}{\hfill\text{(Q.e.d.)}\end{Proof}}

\DeclareMathOperator{\gr}{Gr}
\DeclareMathOperator{\hgr}{\mathfrak{Gr}}
\DeclareMathOperator{\Id}{Id}
\DeclareMathOperator{\Pic}{Pic}
\DeclareMathOperator{\rank}{rank}
\DeclareMathOperator{\spec}{Spec}

\newcommand{\wvarphi}{\widetilde{\varphi}}

\newcommand{\B}{\mathbb{B}}
\newcommand{\C}{\mathbb{C}}
\newcommand{\K}{\mathbb{K}}
\newcommand{\bP}{\mathbb{P}}
\newcommand{\Q}{\mathbb{Q}}
\newcommand{\R}{\mathbb{R}}
\newcommand{\Z}{\mathbb{Z}}

\newcommand{\cE}{\mathcal{E}}
\newcommand{\cF}{\mathcal{F}}
\newcommand{\cK}{\mathcal{K}}
\newcommand{\cO}{\mathcal{O}}
\newcommand{\cQ}{\mathcal{Q}}

\newcommand{\fE}{\mathfrak{E}}
\newcommand{\fF}{\mathfrak{F}}
\newcommand{\fQ}{\mathfrak{Q}}
\newcommand{\fS}{\mathfrak{S}}

\newcommand{\rA}{\mathrm{A}}
\newcommand{\rN}{\mathrm{N}}

\newcommand{\uzero}{\underline{0}}

\begin{document}

\begin{abstract}
Considering the so-called Simpson system on smooth projective varieties, defined over an algebraically closed field of characteristic $0$, whose canonical bundle is ample, I give another proof the stability of this Higgs bundle, from which follows another proof of the Guggenheimer-Yau inequality. Where the equality holds, I prove that the discriminant class of the Simpson system vanishes and this Higgs bundle is curve semistable. This result follows from the study of the relations between ampleness and numerically nefness for Higgs bundles which ``feel'' the Higgs field and (semi)stability. Moreover, I obtain another proof of algebraic hyperbolicity of these varieties which furnishes a lower bound on a real positive constant related to this property; to best of my knowledge, this is the first and unique result of this type.
\end{abstract}

\maketitle

\markboth{Curve semistable Higgs bundles and smooth projective varieties whose canonical bundle is ample}{Armando Capasso}

\section*{Introduction}

\noindent Let $X$ be a smooth, complex, projective variety of dimension $n\geq2$. Recall that ${c_k(X)=(-1)^kc_k\left(\Omega_X^1\right)}$, where $\Omega^1_X$ is the \emph{cotangent bundle of} $X$; under the assumption that $K_X$ (the \emph{canonical line bundle}) is \emph{ample}, Yau proved the inequality
\begin{equation}\label{eqGY}
GY(X)\stackrel{def.}{=}(-1)^n\int_X\left[c_2(X)-\frac{n}{2(n+1)}c_1(X)^2\right]\cdot c_1(X)^{n-2}\geq0,
\end{equation}
and the equality holds if and only if $X$ is covered by the complex ball $\B^n$ holomorphically \cite[Theorem 4 and Remark (iii)]{Y:S-T}. Moreover, assuming $n=2$ and that $X$ saturates \eqref{eqGY}, Miyaoka has proved that $K_X$ and $\Omega^1_X$ are both ample \cite[Corollary at page 294]{M:Y:3}. On another hand, under the hypothesis that $K_X$ is big and numerically effective, Miyaoka proved the same inequality for $n=2$ \cite[Theorem 4]{M:Y:2}.\medskip

\noindent Simpson introduced an opportune \emph{Higgs bundle} $\fS=(S,\varphi)$ (see \eqref{eqSS}) over $X$, and he proved the stability of $\fS$ under opportune hypothesis on $X$ (see \cite[Proposition 9.9]{S:CT}). This result has been extended by Greb, Kebekus, Peternell and Taji to opportune singular, complex, projective varieties (\cite[Example 5.3 and Corollary 7.2]{GKPT}). In particular, from all this it follows a generalization of \eqref{eqGY} to this general setting (\cite[Theorem 1.1]{GKPT}). Indeed, by stability of $\fS$ with respect to a polarization $L$ of $X$ (\cite[Proposition 9.8]{S:CT}):
\begin{displaymath}
\int_X\left[c_2(X)-\frac{n}{2(n+1)}c_1(X)^2\right]\cdot L^{n-2}=\int_X\frac{1}{2(n+1)}c_2\left(S\otimes S^{\vee}\right)\cdot L^{n-2}\geq0;
\end{displaymath}
if $L$ is the canonical bundle, one recovers the inequality \eqref{eqGY}. The characteristic class\newline
$\displaystyle\Delta(S)=\frac{1}{2(n+1)}c_2\left(S\otimes S^{\vee}\right)\in H^2(X,\Q)$ is called \emph{discriminant of $S$}.\medskip
%
%\noindent Simpson, using an opportune \emph{Higgs bundle} $\fS=(S,\varphi)$ (see \eqref{eqSS}), generalized Yau's result to compact K\"ahler manifolds $(M,\omega)$ \cite[Propositions 9.8 and 9.9]{S:CT}. This is accomplished proving the $\omega$-\emph{stability} of that Higgs bundle which implies
%\begin{displaymath}
%GY(M)=\int_M\left[c_2(M)-\frac{n}{2(n+1)}c_1(M)^2\right]\cdot\omega^{n-2}=\int_M\frac{1}{2(n+1)}c_2\left(S\otimes S^{\vee}\right)\cdot\omega^{n-2}\geq0.
%\end{displaymath}
%The characteristic class $\displaystyle\Delta(S)=\frac{1}{2(n+1)}c_2\left(S\otimes S^{\vee}\right)\in H^2(M,\Q)$ is called \emph{discriminant of $S$}.\medskip

\noindent From now on, let $X$ be defined over an algebraically closed field of characteristic $0$. Under the hypothesis that $n\geq2$, $K_X$ is big and numerically effective, Miyaoka proved an inequality similar to \eqref{eqGY} \cite[Proposition 7.1]{M:Y:1}.\medskip

\noindent Assuming the ampleness of $K_X$, I give another proof the stability of $\fS$ with respect to $K_X$ (Theorem \ref{th3.1}) from which it follows another proof of the inequality \eqref{eqGY} (Theorem \ref{th3.3}). In the extremal case, \emph{i.e.} where $\Delta(S)=0$, I prove that $\fS$ is \emph{curve semistable} (Definition \ref{def1.1} and Corollary \ref{cor3.1}.\ref{cor3.1.a}) hence it is semistable with respect to any polarization of $X$ (Corollary \ref{cor3.1}.\ref{cor3.1.c}) and the equality $\displaystyle c_2(X)-\frac{n}{2(n+1)}c_1(X)^2=0\in\rA^2(X)_{\Q}$ holds (Corollary \ref{cor3.1}.\ref{cor3.1.b}).\medskip

\noindent The study of curve semistable Higgs bundles, in the complex setting, started by Bruzzo and Hern\'andez Ruip\'erez in \cite{B:HR} and it had been refined by Bruzzo and Gran\~na Otero in \cite{B:GO:1,B:GO:2}. In \cite{A:C} I proved that the study curve semistability of Higgs bundle over smooth projective varieties, defined over an algebraically closed field of characteristic $0$, may be reduced to complex setting by suitable base changes (Corollary 1.11 \emph{ibidem}).\medskip

\noindent On another hand, in \cite{B:GO:1,B:GO:2} the authors defined notions of ampleness and \emph{numerical effectiveness} for Higgs bundles (Definition \ref{def2.1}, see also Remark \ref{rem2.1}.\ref{rem2.1.a}), using a generalization of the Grassmann bundles of a vector bundle, called the \emph{Higgs Grassmannian schemes}, introduced in \cite{B:HR}. The basic idea is to formulate positivity notions that ``feel'' the Higgs field, and reduce to the classical ones where the Higgs field is zero. In \cite{B:C:GO} Bruzzo, myself and Gra\~na Otero proved the basic properties of ample Higgs bundles (sections 2 and 3 \emph{ibidem}) and applied these results to complex surfaces of general type which saturates \eqref{eqGY} (section 4 \emph{ibidem}).\medskip

\noindent In this paper, I prove a new criterion about the \emph{H-ampleness} (Theorem \ref{th2.3}) which extends the list of analogous criteria proved in the previous paper.\medskip

\noindent From Section \ref{SSMinVarGenType} on, assuming also that $X$ saturates \eqref{eqGY}, from curve semistability of Simpson system $\fS$ it follows another result on $\fS$ (Theorem \ref{th4.1}, cfr. \cite[Proposition 5 at page 293]{M:Y:3}) which imposes some restriction on the curves over $X$ (Corollary \ref{cor4.1}). Since $K_X$ is ample, it follows also the H-ampleness of $\fS$ (Lemma \ref{lem3.1}) applying \cite[Corollary 3.2]{B:C:GO}. From all this, it follows a proof of the ampleness of $\Omega^1_X$ (Corollary \ref{cor3.3}).\medskip

\noindent Moreover, in the complex setting, the stability of $\fS$ allows me to give a new proof of the fact that $X$ is uniformized by the complex ball $\B^n$ if $X$ saturates \eqref{eqGY} (Theorem \ref{th5.1}), even if there are general results in opportune singular, complex, projective settings (\cite[Theorems 1.2 and 1.3]{GKPT}). Instead, the curve semistability of $\fS$ allows me to give a new proof of the \emph{algebraic hyperbolicity} of $X$ (Definition \ref{def5.1} and Theorem \ref{th3.5}). The advantage of this new proof is to obtain a lower bound for a real positive constant which depends on $X$ and $K_X$.\medskip

\noindent Finally, assuming the ampleness of the cotangent bundle $\Omega^1_X$ of $X$ instead of the ampleness of $K_X$, I prove the H-ampleness of $\fS$ (Lemma \ref{lem3.2}); this last result is relied on the proof of Theorem \ref{th3.1}. In particular, it happens that $\Omega^1_X$ is ample if and only if $\fS$ is H-ample (Proposition \ref{prop3.1}).\medskip

\noindent \textbf{Notations and conventions.} $\K$ is an algebraically closed field of characteristic $0$, unless otherwise indicated. By a projective variety $X$ I mean a projective integral scheme over $\K$ of dimension $n\geq1$ and of finite type. If $n\in\{1,2\}$ I shall write projective curve or projective surface, respectively. Whenever I consider a morphism $f\colon C\to X$, I understand that $C$ is an irreducible smooth projective curve.\medskip

\noindent{\bf Acknowledgment.} I am very grateful to my Ph.D. advisors Ugo Bruzzo and Beatriz Gra\~na Otero for their help, their energy and their support. I thank SISSA for the hospitality while part of this work was done. I thank Angelo Felice Lopez for pointing out some oversights to me in the references.

\section{Curve semistable Higgs bundles}

\noindent Let $X$ be a smooth scheme over $\K$, let $\Omega^1_X$ be the cotangent bundle of $X$.
\begin{definition}
A \emph{Higgs sheaf} $\fE$ on $X$ is a pair $(\cE,\varphi)$ where $\cE$ is an $\cO_X$-coherent sheaf equipped with a morphism $\varphi\colon\cE\to\cE\otimes\Omega^1_X$ called \emph{Higgs field} such that the composition
\begin{displaymath}
\varphi\wedge\varphi\colon\cE\xrightarrow{\varphi}\cE\otimes\Omega^1_X\xrightarrow{\varphi\otimes\Id}\cE\otimes\Omega^1_X\otimes\Omega^1_X\to\cE\otimes\Omega^2_X
\end{displaymath}
vanishes. A Higgs subsheaf of $\fE$ is a $\varphi$-\emph{invariant subsheaf} $\cF$ of $\cE$, \emph{i.e.} $\varphi(\cF)\subseteq\cF\otimes\Omega_X^1$. A \emph{Higgs quotient} of $\fE$ is a quotient sheaf of $\cE$ such that the corresponding kernel is $\varphi$-invariant. A \emph{Higgs bundle} is a Higgs sheaf whose underlying coherent sheaf is locally free.
\end{definition}
\noindent Let $H$ be a polarization of $X$ and let $\fE=(\cE,\varphi)$ be a torsion-free Higgs sheaf on $X$, if not otherwise indicated.
One defines the \emph{slope} of $\fE$ as $\displaystyle\mu(\fE)=\frac{1}{\rank(\cE)}\int_Xc_1(\det(\cE))\cdot H^{n-1}\in\Q$.
\begin{definition}
$\fE$ is $H$-(\emph{semi})\emph{stable} if $\mu(\fF)\stackrel[(\leq)]{\textstyle<}{}\mu(\fE)$ for every Higgs subsheaf $\fF$ of $\fE$ with ${0<\rank(\fF)<\rank(\fE)}$. Or $\fE$ is $H$-(\emph{semi})\emph{stable} if and only if $\mu(\fE)\stackrel[(\leq)]{\textstyle<}{}\mu(\fQ)$ for every torsion-free Higgs quotient sheaf $\fQ$ of $\fE$ with $0<\rank(\fQ)<\rank(\fE)$, equivalently. In the other eventuality, $\fE$ is \emph{unstable}.
\end{definition}
\begin{definition}\label{def1.1}
$\fE$ is \emph{curve semistable} if $f^{\ast}\fE$ is semistable for any $f\colon C\to X$.
\end{definition}
\noindent For simplicity, I shall skip any reference to the fixed polarization $H$ of $X$ if there is no confusion.\medskip

\noindent I consider the characteristic class
\begin{displaymath}
\Delta(\cE)=\frac{1}{2r}c_2\left(\cE\otimes\cE^{\vee}\right)=c_2(\cE)-\frac{r-1}{2r}c_1(\cE)^2\in\rA^2(X)\otimes_{\Z}\Q=\rA^2(X)_{\Q},
\end{displaymath}
called the \emph{discriminant} of $\cE$ (here $r=\rank(\cE)$). 
Here $\rA_k(X)$ is the Abelian group of $k$-cycles on $X$ modulo rational equivalence and $\rA^k(X)=\rA_{n-k}(X)$.
\begin{theorem}[{\cite[Theorem 7]{L:A}}]\label{th1.1}
Let $\fE=(\cE,\varphi)$ be a semistable Higgs sheaf on $X$ with respect to a polarization $H$. Then
\begin{displaymath}
\int_X\Delta(\cE)\cdot H^{n-2}\geq0.
\end{displaymath}
\end{theorem}
\begin{remark}
If one assumes $\K=\C$, the previous theorem is \cite[Proposition 3.4]{S:CT}.
\end{remark}
\noindent About the ``extremal'' case, it subsists the following theorem.
\begin{theorem}[{see \cite[Theorems 1.2 and 1.3]{B:HR} and \cite[Proposition 3.2]{L:LG}}]\label{th1.2}
Let $\fE=(E,\varphi)$ be a Higgs bundle over $X$.
\begin{enumerate}[a)]
\item\label{th1.2.a} If $\fE$ is semistable with respect to some polarization $H$ and $\displaystyle\int_X\Delta(E)\cdot H^{n-2}=0$. Then $\fE$ is curve semistable.
\item\label{th1.2.b} If $\fE$ is curve semistable. Then $\fE$ is semistable with respect to some polarization $H$.
\end{enumerate} 
\end{theorem}
\begin{remarks}\label{rem1.1}
\,\begin{enumerate}[a)]
\item\label{rem1.1.a} In statement \ref{th1.2.a} of the previous Theorem, one may write $\Delta(E)=0$ instead of $\displaystyle{\int_X\Delta(E)\cdot H^{n-2}=0}$. Indeed, by replacing $E$ with $E\otimes E^{\vee}$ if needed, one may assume that $c_1(E)=0$. By hypothesis, this allows one to apply \cite[Corollary 6]{L:A}, so that $c_k(E)=0$ for all $k>0$, and then $\Delta(E)=0$.
\item Statement \ref{th1.2.b} of the previous theorem follows from the ``easy'' direction of \cite[Theorem 6.1]{M:R}. It is an open problem either to prove or to disprove the vanishing of the discriminant class of a curve semistable Higgs bundle.
\end{enumerate}
\end{remarks}

\section{H-ample and H-nef Higgs bundles}

\noindent Let $\fE=(E,\varphi)$ be a rank $r\geq2$ Higgs bundle over a smooth projective variety $X$, and let ${s\in\{1,\dotsc,r-1\}}$ be an integer number. Let $p_s\colon\gr_s(E)\to X$ be the \emph{Grassmann bundle} parametrizing rank $s$ locally free quotients of $E$ (see \cite{F:W}). Consider the short exact sequence of vector bundles over $\gr_s(E)$ 
\begin{displaymath}
\xymatrix{
0\ar[r] & S_{r-s,E}\ar[r]^(.55){\eta} & p_s^{\ast}E\ar[r]^(.45){\epsilon} & Q_{s,E}\ar[r] & 0
},
\end{displaymath}
where $S_{r-s,E}$ is the \emph{universal rank $r-s$ subbundle} and $Q_{s,E}$ is the \emph{universal rank $s$ quotient bundle of} $p_s^{\ast}E$, respectively. One defines the closed subschemes $\hgr_s(\fE)\subseteq\gr_s(E)$ (the $s$\emph{-th Higgs-Grassmann schemes of} $\fE$) as the zero loci of the composite morphisms
\begin{displaymath}
(\epsilon\otimes\Id)\circ p_s^{\ast}\varphi\circ\eta\colon S_{r-s,E}\to Q_{s,E}\otimes p_s^{\ast}\Omega_X^1.
\end{displaymath}
The restriction of the previous sequence to $\hgr_s(\fE)$ yields a universal short exact sequence
\begin{displaymath}
\xymatrix{
0\ar[r] & \mathfrak{S}_{r-s,\fE}\ar[r]^(.55){\psi} & \rho_s^{\ast}\fE\ar[r]^(.45){\eta} & \fQ_{s,\fE}\ar[r] & 0,
}
\end{displaymath}
where $\fQ_{s,\fE}=Q_{s,E|\hgr_s(\fE)}$ is equipped with the quotient Higgs field induced by $\rho_s^{\ast}\varphi$, where ${\rho_s=p_{s|\hgr_s(\fE)}}$.
The scheme $\hgr_s(\fE)$ enjoys the usual universal property: for a morphism of varieties $f\colon Y\to X$, the morphism $g\colon Y \to \gr_s(E)$ given by a rank $s$ quotient $Q$ of $f^{\ast} E$ factors through $\hgr_s(\fE)$ if and only if $\varphi$ induces a Higgs field on $Q$.
\begin{definition}[{see \cite[Definition 2.3]{B:GO:1}}]\label{def2.1}
A Higgs bundle $\fE=(E,\varphi)$ of rank one is said to be \emph{Higgs ample}\textbackslash\emph{Higgs numerically effective} (\emph{H-ample}\textbackslash\emph{H-nef}, for short) if $E$ is ample\textbackslash numerically effective in the usual sense. If $\rank(\fE)\geq 2$, one defines H-ampleness\textbackslash H-nefness inductively by requiring that
\begin{enumerate}[a)]
\item all Higgs bundles $\fQ_{s,\fE}$ are H-ample\textbackslash H-nef for all $s$, and
\item the determinant line bundle $\det(E)$ is ample\textbackslash numerically effective.
\end{enumerate}
\end{definition}
\begin{remarks}\label{rem2.1}
\,\begin{enumerate}[a)]
\item\label{rem2.1.a} Following \cite{L:RK}, a line bundle $L$ over a proper scheme $Z$ is \emph{numerically effective} (\emph{nef}, for short) if $\displaystyle\int_Cc_1(L)\geq0$ for every irreducible curve $C\subseteq Z$.
\item In the previous definition, the condition on the determinant cannot be omitted as \cite[Example 2.5]{B:GO:1} shows.
\item\label{rem2.1.c} Let $\fE=(E,\varphi)$ be a Higgs bundle over $X$. Where $\varphi=0$, $\fE$ is H-ample\textbackslash H-nef if and only if $E$ is ample\textbackslash numerically effective as ordinary vector bundle.
\item\label{rem2.1.d} The recursive condition in this definition can be recast as follows. Let ${1\leq s_1<s_2<\dotsc<s_k<r}$ and let $\fQ_{\left(s_1 \cdots,s_k\right),\fE}$ be the rank $s_1$ universal Higgs quotient bundle over $\hgr_{s_1}\left(\fQ_{\left(s_2,\dotsc,s_k\right),\fE}\right)$, obtained by taking the successive universal Higgs quotient bundles of $\fE$ of rank $s_k$, then $s_{k-1}$, all the way to rank $s_1$. The H-ampleness condition for $\fE$ is equivalent to requiring that all line bundles $\det(\fE)$ and $\det(\fQ_{\left(s_1,\cdots,s_k\right),\fE})$ are ample. \hfill{$\Diamond$}
\end{enumerate}
\end{remarks}
\begin{propositions}\label{prop2.1}
\,
\begin{enumerate}[a)]
\item\label{prop2.1.a} Let $f\colon Y\to X$ be a finite morphism of smooth projective varieties. If $\fE$ is H-ample then $f^{\ast}\fE$ is H-ample. Moreover, if $f$ is also surjective and $f^{\ast}\fE$ is H-ample then $\fE$ is H-ample (\cite[Proposition 2.6.(1)]{B:C:GO}).
\item\label{prop2.1.b} Let $\fE$ be H-ample then every Higgs quotient bundle of $\fE$ is H-ample (\cite[Proposition 2.6.(2)]{B:C:GO}).
%\item\label{prop2.1.c} Let $\fE=(E,\varphi)$ be curve semistable and let $c_1(E)$ be ample. Then $\fE$ is H-ample (\cite[Proposition 3.11]{B:C:GO}).
\end{enumerate}
\end{propositions}
\noindent I recall that every Higgs bundle $\fE$ has a \emph{Harder-Narasimhan filtration} (\emph{HN-filtration}, for short)
\begin{equation}\label{eqHNfiltr}
{0}=\fE_0\subsetneqq\fE_1\subsetneqq\dotsc\subsetneqq\fE_{m-1}\subsetneqq\fE_m=\fE
\end{equation}
and the slopes of the successive quotients $\fQ_i=\fE_i/\fE_{i-1}$ satisfy the condition
\begin{displaymath}
\mu_{\max}(\fE)\stackrel{def.}{=}\mu(\fQ_1)>\dotsc>\mu(\fQ_m)\stackrel{def.}{=}\mu_{\min}(\fE).
\end{displaymath}
For every Higgs quotient $\fQ$ of $\fE$ one has $\mu(\fQ)\geq\mu_{\min}(\fE)$.
\begin{theorem}[{\cite[Theorem 3.1]{B:C:GO}}]\label{th2.1}
Let $\fE=(E,\varphi)$ be a Higgs bundle over $X$. Fix an ample class $h\in\rN^1(X)$. Then $\fE$ is H-ample if and only if
\begin{enumerate}[a)]
\item the line bundle $\det(E)$ is ample;
\item there exists $\delta\in\R_{>0}$ such that for every finite morphism $f\colon C\to X$, the inequality
\begin{displaymath}
\mu_{\min}\left(f^{\ast}\fE\right)\geq\delta\int_Cf^{\ast}h
\end{displaymath}
holds, where $\mu_{\min}\left(f^{\ast}\fE\right)$ is defined via the relevant HN-filtration \eqref{eqHNfiltr}.
\end{enumerate}
\end{theorem}
\noindent Moreover, it subsists the following theorem.
\begin{theorem}\label{th2.3}
Let $\fE=(E,\varphi)$ be a Higgs bundle over $X$. $\fE$ is H-ample if and only if
\begin{enumerate}[a)]
\item the line bundle $\det(E)$ is ample;
\item\label{th2.3.b} for each positive-dimensional subvariety $Y$ of $X$, the Higgs quotient bundles $\fQ$ of $\fE_{\vert Y}$ are H-ample.
\end{enumerate}
\end{theorem}
\begin{prf}
The ``only if'' part follows by definition and Proposition \ref{prop2.1}.\ref{prop2.1.b}.\smallskip

\noindent Let assume that $\det(E)$ is ample and condition \ref{th2.3.b} holds. The H-ampleness of $\fE$ is equivalent to the ampleness of a collection of line bundles $L_Z$, each on an iterated Higgs Grassmannian, which we denote generically by $Z$, with projection $\rho_Z\colon Z\to X$ (these line bundles are obtained by successively taking the universal Higgs quotient until one reaches the rank one quotient bundles, see Remark \ref{rem2.1}.\ref{rem2.1.d}). Let $\zeta_Z\colon\rho_Z^{\ast}\fE\to L_Z$ be the quotient morphism, let $g\colon W\hookrightarrow Z$ be a positive-dimensional subvariety of $Z$, and let $f=\rho_Z\circ g$. We have a Higgs quotient $f^{\ast}\fE\to\fQ$, where $\fQ=g^{\ast}L_Z$. By condtion \ref{th2.3.b} and by \emph{Nakai-Moishezon-Kleiman Criterion}
\begin{displaymath}
\int_{g(W)}c_1\left(L_Z\right)^{\dim g(W)}=\int_Wc_1(\fQ)^{\dim W}>0,
\end{displaymath}
so that $L_Z$ is ample for every choice of $Z$. As a consequence, $\fE$ is H-ample.
\end{prf}
\noindent Also \emph{H-nef} Higgs bundles satisfy analogous properties of H-ample Higgs bundles.
\begin{propositions}\label{prop2.2}
\,\begin{enumerate}[a)]
\item\label{prop2.2.a} Let $f\colon Y\to X$ be a morphism of smooth projective varieties. Then $f^{\ast}\fE$ is H-nef (\cite[Proposition 2.6.(ii)]{B:GO:1}). If $f$ is also surjective and $f^{\ast}\fE$ is H-nef then $\fE$ is H-nef (\cite[Lemma 3.4]{B:B:G}).
\item\label{prop2.2.b} Let $\fE$ be H-nef then every Higgs quotient bundle of $\fE$ is H-nef (\cite[Lemma 3.5]{B:B:G}).
\end{enumerate}
\end{propositions}
\begin{theorem}[{\cite[Theorem 3.4]{B:C:GO}}]\label{th2.2}
Let $\fE$ be a Higgs bundle over $X$. Then $\fE$ is H-nef if and only if $\mu_{\min}\left(f^{\ast}\fE\right)\geq0$ for all morphisms $f\colon C\to X$, where $\mu_{\min}\left(f^{\ast}\fE\right)$ is defined via the relevant HN-filtration \eqref{eqHNfiltr}.
\end{theorem}

\section{The Simpson system on smooth projective varieties\\ whose canonical bundle is ample}\label{SSMinVarGenType}

\noindent From now on, let $X$ be a smooth projective variety of dimension $n\geq2$ with $K_X$ (the \emph{canonical line bundle}) ample, unless otherwise indicated. Let $\fS=(S,\varphi)$ be the \emph{Simpson system}, where $S=\Omega^1_X\oplus\cO_X$ and
\begin{equation}\label{eqSS}
\varphi=\begin{pmatrix}
0 & 0\\
\Id & 0
\end{pmatrix},\,\Id\in\hom\left(\Omega^1_X,\Omega^1_X\right).
\end{equation}
Actually the Simpson system is $K_X$-stable.
\begin{theorem}
The cotangent bundle $\Omega^1_X$ is $K_X$-semistable.
\end{theorem}
\begin{prf}
Repeating the reasoning of \cite[Theorem 1.7]{A:C}, one can assume $\K\subseteq\C$. Consider the following Cartesian diagram
\begin{displaymath}
\xymatrix{
X_{\C}\ar[r]^{f}\ar[d] & X\ar[d]\\
\spec(\C)\ar[r] & \spec(\K)
}
\end{displaymath}
since $X_{\C}$ is a smooth projective variety of dimension $n$ (\cite[Lemma A.3]{A:C}), $f^{\ast}K_X\cong K_{X_{\C}}$ and $f^{\ast}\Omega^1_X\cong\Omega^1_{X_{\C}}$ ( \cite[\href{https://stacks.math.columbia.edu/tag/01V0}{tags 01V0} and \href{https://stacks.math.columbia.edu/tag/0D2P}{0D2P}]{TSP}). $K_{X_{\C}}$ is ample hence $\Omega^1_{X_{\C}}$ is $K_{X_{\C}}$-semistable (see \cite[Theorem 1]{T:H}, for example). From all this, the claim follows by \cite[Lemma 1.10]{A:C}.
\end{prf}
\begin{theorem}\label{th3.1}
The Higgs bundle $\fS$ is $K_X$-stable.
\end{theorem}
\begin{prf}
By Bertini's Theorem (cfr. \cite[Corollary III.10.9 and Exercise III.11.3]{H:RC}), for $m\gg1$ there exist general, ample, smooth divisors $D_1,\dotsc,D_{n-1}\in\left|mK_X\right|$ such that $Y_p=D_1\cap\dotsc\cap D_p$ are smooth projective subvarieties of $X$ for any $p\in\{1,\dotsc,n-1\}$\footnote{
Moreover, these smooth projective varieties are all minimal and of general type. Indeed, by Adjunction Formula $K_{Y_1}=\left(K_X+Y_1\right)_{\vert Y_1}$ (\cite[Exercise 21.5.B]{FOAG}) and this is an ample line bundle (\cite[Proposition 1.2.13 and Corollary 1.4.10]{L:RK}). Iterating these reasoning for all $Y_p$ one has the claim.
}; one puts $C=Y_{n-1}$. Since $\Omega^1_X$ is $K_X$-semistable, by \cite[Theorem 6.1]{M:R} $\Omega^1_{X\vert C}$ is semistable.\smallskip

\noindent The Higgs field $\psi$ induced by $\varphi_{\vert C}$ on the Higgs quotient $\Omega^1_{X\vert C}$ vanishes. Indeed, reasoning on the stalks, one has
\begin{gather*}
\forall x\in C,\widetilde{\omega}\in\Omega^1_{X\vert C},\,\epsilon_x^{-1}\left(\widetilde{\omega}\right)=\left\{(s,\omega)\in\fS_x\mid\pi(\omega)=\widetilde{\omega}\right\},\\
\varphi_{\vert C,x}\left(\epsilon_x^{-1}\left(\widetilde{\omega}\right)\right)=\left\{\widetilde{\omega}\right\}\subseteq\Omega^1_{C,x}=\ker(\epsilon\otimes\Id)_x
\end{gather*}
where $\pi\colon\Omega^1_{X\vert C}\to\Omega^1_C$ is the canonical projection. Thus $\psi$ vanishes.\medskip

\noindent One has the following inequality
\begin{displaymath}
0<\mu\left(S_{\vert C}\right)=\frac{1}{n+1}\int_Cc_1\left(S_{\vert C}\right)<\frac{1}{n}\int_Cc_1\left(\Omega^1_{X\vert C}\right)=\mu\left(\Omega^1_{X\vert C}\right).
\end{displaymath}
Let $\varpi\colon\fS_{\vert C}\to\fQ=\left(\cQ,\wvarphi\right)$ be a rank $s$ Higgs quotient bundle of $\fS$ and let $\cK=\ker(\varpi)$; $\cK$ is locally free because it is a kernel of an epimorphism of locally free sheaves on a Noetherian scheme. One sets
\begin{displaymath}
\cK_1=\cK\cap\Omega^1_{X\vert C},\cK_2=\cK\cap\cO_C.
\end{displaymath}
\begin{enumerate}[1)]
\item Let $\cK_2=\uzero_C$ then $\cK\subseteq\Omega^1_{X\vert C}$. This cannot happen, indeed, by assumption ${\varphi_{\vert C}(\cK)\subseteq\cK\otimes\Omega^1_C}$ but by definition $\varphi_{\vert C}(\cK)=\uzero_C\oplus\pi(\cK)\subseteq\left(\Omega^1_{X\vert C}\otimes\Omega^1_C\right)\oplus\Omega^1_C$, \emph{i.e.} $\cK$ cannot be a Higgs subsheaf of $\fS_{\vert C}$. In other words, $\Omega^1_{X\vert C}$ does not contain any Higgs subsheaves of $\fS_{\vert C}$.
\item Let $\cK_2\neq\uzero_C$ then $\cK_2=\cO_C$. Indeed, consider the following commutative diagram
\begin{displaymath}
\xymatrix{
 & 0\ar[d] & 0\ar[d]\\
0\ar[r] & \cK_2\ar@{^{(}->}[d]\ar@{^{(}->}[r] & \cO_C\ar@{^{(}->}[d]\\
0\ar[r] & \cK\ar@{->>}[d]\ar@{^{(}->}[r] & S_{\vert C}\ar@{->>}[d]\ar@{->>}[r] & \cQ\ar@{->>}[d]^q\ar[r] & 0\\
0\ar[r] & \cK_1\ar[d]\ar@{^{(}->}[r] & \Omega^1_{X\vert C}\ar[d]\ar@{->>}[r] & \cQ_0\ar[d]\ar[r] & 0\\
 & 0 & 0 & 0
}
\end{displaymath}
where the columns and the rows are short exact sequence of sheaves. By the universal property of cokernels of morphisms, there exists a unique morphism $q\colon\cQ\twoheadrightarrow\cQ_0$ which makes commutative the diagram, and it is also an epimorphism. Computing the ranks:
\begin{gather*}
\rank\left(S_{\vert C}\right)=n+1,\rank(\cQ)=s\Rightarrow\rank(\cK)=n-s+1\\
\rank\left(\cK_2\right)=1,\rank(\cK)=n-s+1\Rightarrow\rank\left(\cK_1\right)=n-s\\
\rank\left(\Omega^1_{X\vert C}\right)=n,\rank\left(\cK_1\right)=n-s\Rightarrow\rank\left(\cQ_0\right)=s\\
\rank(\cQ)=\rank\left(\cQ_0\right)=s\Rightarrow\rank\left(\ker(q)\right)=0;
\end{gather*}
on the other hand, $\ker(q)$ is locally free, because it is a torsion-free coherent sheaf on a smooth curve. Thus $\ker(q)=\uzero_C$ and $\cQ\cong\cQ_0$. Applying the Snake Lemma, one has the long exact sequence
\begin{displaymath}
\xymatrix{
0\ar[r] & \cK_2\ar@{^{(}->}[r] & \cO_C\ar[r] & \ker(q)=\uzero_C
}
\end{displaymath}
\emph{i.e.} $\cK_2=\cO_C$.
\end{enumerate}
From all this, it turns out that $\fQ$ is a Higgs quotient bundle of $\left(\Omega^1_{X\vert C},0\right)$. Since $\Omega^1_{X\vert C}$ is semistable, $\mu\left(S_{\vert C}\right)<\mu\left(\Omega^1_{X\vert C}\right)\leq\mu(Q)$ hence the claim follows.
\end{prf}
\noindent Recall
\begin{displaymath}
GY(X)=(-1)^n\int_X\left[c_2(X)-\frac{n}{2(n+1)}c_1(X)^2\right]\cdot K_X^{n-2};
\end{displaymath}
where $c_k(X)\stackrel{def.}{=}(-1)^kc_k\left(\Omega^1_X\right)$ for any $k$. It is known that $GY(X)\geq0$ (cfr. \cite[Proposition 7.1]{M:Y:1}) which is called \emph{Guggenheimer-Yau Inequality} (see \cite[Remark (iii)]{Y:S-T}). By the previous theorem it follows another proof of this inequality.
\begin{theorem}[{cfr. \cite[Proposition 7.1]{M:Y:1}}]\label{th3.3}
The Guggenheimer-Yau inequality ${GY(X)\geq0}$ holds.
\end{theorem}
\begin{prf}
It is enough to note that 
\begin{displaymath}
GY(X)=\int_X\Delta(S)\cdot c_1\left(K_X\right)^{n-2}
\end{displaymath}
and the claim follows from Theorem \ref{th1.1}.
\end{prf}

\subsection{Smooth projective varieties of general type such that $GY(X)=0$}

\noindent In this subsection, I assume $GY(X)=0$ unless otherwise indicated.\medskip

\noindent By Theorem \ref{th1.2} and Remark \ref{rem1.1}.\ref{rem1.1.a} the following corollary holds.
\begin{corollary}\label{cor3.1}
\,\begin{enumerate}[a)]
\item\label{cor3.1.a} The Higgs bundle $\fS$ is curve semistable.
\item\label{cor3.1.b} $\displaystyle\Delta(S)=c_2(X)-\frac{n}{2(n+1)}c_1(X)^2=0\in\rA^2(X)_{\Q}$.
\item\label{cor3.1.c} The Higgs bundle $\fS$ is semistable with respect to any polarization of $X$.
\end{enumerate}
\end{corollary}
\begin{lemma}\label{lem3.1}
The Higgs bundle $\fS$ is H-ample.
\end{lemma}
\begin{prf}
Fix an ample class $h$. By the H-ampleness criterion (Theorem \ref{th2.1}) a Higgs bundle $\fE$ is H-ample if and only if its determinant is ample and there exists a $\delta\in\R_{>0}$ such that
\begin{displaymath}
\mu_{\min}(f^{\ast}\fE)\geq\delta\int_C f^{\ast}h
\end{displaymath}
for all $f\colon C\to X$. One takes $h=K_X$ and $\displaystyle\delta=\frac{1}{n+1}$. Since $\fS$ is curve semistable (Corollary \ref{cor3.1}.\ref{cor3.1.a})
\begin{displaymath}
\mu_{\min}\left(f^{\ast}\fS\right)=\mu\left(f^{\ast}\fS\right)=\frac{1}{n+1}\int_Cf^{\ast}c_1(S)=\delta\int_Cf^{\ast}K_X.
\end{displaymath}
\end{prf}
\begin{remark}
$S$, as an ordinary vector bundle, is neither stable nor ample. Indeed, $\cO_X$ is a quotient of $S$ and $\mu(S)>\mu(\cO_X)=0$.
\end{remark}
\noindent By the proof of Theorem \ref{th3.1}, applying Proposition \ref{prop2.1}.\ref{prop2.1.b}, one has the following corollary.
\begin{corollary}\label{cor3.3}
The cotangent bundle $\Omega^1_X$ is ample.
\end{corollary}

\subsection{Smooth projective varieties of general type whose cotangent bundle is ample}

\noindent In this subsection, I assume $\Omega^1_X$ ample. Also under this hypothesis, Lemma \ref{lem3.1} holds.
\begin{lemma}\label{lem3.2}
The Higgs bundle $\fS$ is H-ample.
\end{lemma}
\begin{prf}
Since $\det(S)=K_X$, the condition on the determinant line bundle is fulfilled by the assumption. By the proof of Theorem \ref{th3.1}, the Higgs quotient bundles of $\fS$ are ample, so that the claim follows by Theorem \ref{th2.3}.
\end{prf}
\noindent However a smooth projective varieties whose cotangent bundle is ample can not saturate Guggenheimer-Yau inequality.
\begin{example}
Let $Y$ be a smooth complex $3$-fold of general type whose cotangent bundle is ample, for example $Y$ is a quotient of the unit ball of $\C^3$. A smooth hyperplane section $X$ of $Y$ has ample cotangent bundle, hence it is a smooth projective surface of general type (see \cite[Section 3.3]{D:O:1}). Let $aL$ be the class of $X$ in $\rA^1(Y)$; for $a\gg1$ one has $GY(X)>0$.
\end{example}
\noindent In summary, Lemmata \ref{lem3.1}, \ref{lem3.2} and Corollary \ref{cor3.3} prove the following proposition.
\begin{proposition}\label{prop3.1}
Let $X$ be a smooth projective variety whose canonical bundle is ample. Then $\Omega^1_X$ is ample if and only if $\fS$ is H-ample.
\end{proposition}
\begin{example}
Let $C_2$ and $C_2$ be irreducible, smooth, projective curves of genus greater or equal to $2$. Let $X=C_1\times C_2$; this is a smooth projective surface with $K_X$ ample and $\Omega^1_X$ nef but not ample. Then $\fS$ is a H-nef but not H-ample Higgs bundle over $X$ with ample determinant line bundle.
\end{example}

\section{On twisted Simpson system}

\noindent From now on, let $X$ be a smooth projective variety of dimension $n\geq2$ with $K_X$ ample and $GY(X)=0$, unless otherwise indicated.\medskip

\noindent More in general than Lemma \ref{lem3.1}, the following theorem holds.
\begin{theorem}\label{th4.1}
The Higgs bundle $\fS_{\beta}=\fS\left(-\beta K_X\right)$ is:
\begin{enumerate}[a)]
\item\label{th4.1.a} H-nef for every rational number\footnote{
As it is customary, I consider twistings by rational divisors formally, which make sense after pulling back to a (possibly ramified) finite covering of $X$; on the other hand, the properties of being semistable, H-ample, H-nef are invariant under such coverings (see Propositions \ref{prop2.1}.\ref{prop2.1.a} and \ref{prop2.2}.\ref{prop2.2.a}).
} $\displaystyle\beta\leq\frac{1}{n+1}$;
\item\label{th4.1.b} H-ample for every rational number $\displaystyle\beta<\frac{1}{n+1}$.
\end{enumerate}
\end{theorem}
\begin{prf}
By Theorem \ref{th3.1}, after twisting by a line bundle, and by \cite[Lemma 2.3]{B:HR} $\fS_{\beta}$ is curve semistable, so that for every morphism $f\colon C\to X$, the pullback Higgs bundle $f^{\ast}\fS_{\beta}$ is semistable hence
\begin{displaymath}
\mu\left(f^{\ast}\fS_{\beta}\right)=\mu_{\min}\left(f^{\ast}\fS_{\beta}\right)=\frac{1}{n+1}\int_Cf^{\ast}c_1\left(S_{\beta}\right)=\frac{1}{n+1}\int_{\overline{C}}c_1\left(S\left(-\beta K_X\right)_{\vert\overline{C}}\right)=\left(\frac{1}{n+1}-\beta\right)\int_{\overline{C}}K_{X\vert\overline{C}}\geq0 
\end{displaymath}
where $\overline{C}=f(C)$; the last inequality holds as $K_X$ is ample. By the properties of the HN-filtration of $f^{\ast}\fE$, any Higgs quotient of $f^{\ast}\fS_{\beta}$ has non negative degree. Furthermore, by the previous computation
\begin{displaymath}
\int_Cf^{\ast}c_1\left(S_{\beta}\right)=(1-(n+1)\beta)\int_{\overline{C}}K_{X\vert\overline{C}}\geq0.
\end{displaymath}
Thus, by the H-ampleness and H-nefness criteria (Theorems \ref{th2.1} and \ref{th2.2}, respectively), one has the ampleness and nefness of $\det\left(\fS_{\beta}\right)$ in function of $\beta$, so that the claim follows.
\end{prf}
\begin{remark}
For any rational number $\displaystyle0<\beta\leq\frac{1}{n+1}$, $\fS_{\beta}$ is a H-nef Higgs bundle whose underlying vector bundle $S_{\beta}$ is not nef; indeed, $-\beta K_X$ is quotient line bundle of $S_{\beta}$ which is not nef.~\end{remark}
\noindent In consequence, I generalize \cite[Corollary at page 294]{M:Y:3} to higher dimensional case.
\begin{corollary}\label{cor4.1}
The vector bundle $\Omega_{\beta}=\Omega^1_X\left(-\beta K_X\right)$ is:
\begin{enumerate}[a)]
\item nef for every rational number $\displaystyle\beta\leq\frac{1}{n+1}$;
\item ample for every rational number $\displaystyle\beta<\frac{1}{n+1}$.
\end{enumerate}
As consequences for each irreducible projective curve $C$ on $X$ the inequality $\displaystyle2p_a(C)-2\geq\frac{1}{n+1}\int_CK_{X\vert C}$ holds, where $p_a(C)$ is the \emph{arithmetic genus of $C$}. In particular $X$ does not contain curves of arithmetic genus neither $0$ nor $1$.
\end{corollary}
\begin{prf}
$\Omega_{\beta}$ with the zero Higgs field is a Higgs quotient of $\fS_{\beta}$ hence it is H-nef (Proposition \ref{prop2.2}.\ref{prop2.2.b}) and then nef in the usual sense (Remark \ref{rem2.1}.\ref{rem2.1.c}). Let $C\subsetneqq X$ be an irreducible projective curve, let $\left(\widetilde{C},\nu\right)$ be its normalization, and let $\iota\colon\widetilde{C}\xrightarrow{\nu}C\hookrightarrow X$. One has the right exact sequence
\begin{displaymath}
\xymatrix{
\iota^{\ast}\Omega_{\textstyle\frac{1}{n+1}}\ar@{->>}[r] & \displaystyle\Omega^1_{\widetilde{C}}\left(-\frac{1}{n+1}\iota^{\ast}K_X\right)=\cO_{\widetilde{C}}\left(K_{\widetilde{C}}-\frac{1}{n+1}\iota^{\ast}K_X\right)\ar[r] & 0
}
\end{displaymath}
so that $\displaystyle\cO_{\widetilde{C}}\left(K_{\widetilde{C}}-\frac{1}{n+1}\iota^{\ast}K_X\right)$ is nef, \emph{i.e.} its degree is nonnegative, and
\begin{displaymath}
2p_a(C)-2\geq2p_a\left(\widetilde{C}\right)-2=\deg\Omega^1_{\widetilde{C}}\geq\frac{1}{n+1}\int_CK_{X\vert C}>0.
\end{displaymath}
As a consequence $X$ has curves of arithmetic genus neither $0$ nor $1$.
\end{prf}
\noindent On the other hand, the previous inequality is not sharp as the following example proves.
\begin{example}
Let $X$ be a \emph{fake complex projective plane}, \emph{i.e.} $X$ is a smooth projective surface with the same \emph{Betti numbers of} $\bP^2_{\C}$ but is not isomorphic to $\bP^2_{\C}$. Its canonical bundle $K_X$ is ample hence $X$ is a minimal smooth surface of general type and $\displaystyle{\int_X3c_2(X)=\int_Xc_1(X)^2=9}$ and the \emph{Picard number of $X$ is} $1$. Let $\cO_X(1)$ be the ample generator of the torsion-free part of $\Pic(X)$ such that $K_X=\cO_X(3)$ and $c_1\left(\cO_X(1)\right)^2=1$ (cfr. \cite[Section 1.1]{DiC:LF}). Moreover, let $f\colon C\to X$ non-constant then $g(C)\geq3$ by \cite[Lemma 2.2]{K:JH} and \cite[Proposition 2.3]{DiC:LF}. However, for a such $C$ the previous corollary predicts:
\begin{displaymath}
2g(C)-2\geq\frac{1}{2+1}\int_X\cO_X(3)C=\int_X\cO_X(1)C>0\iff g(C)\geq2.
\end{displaymath}
\end{example}
\noindent A little recap: in Lemma \ref{lem3.1}, Theorem \ref{th4.1}, Corollaries \ref{cor3.1} and \ref{cor4.1} $X$ is a smooth projective variety of dimension $n\geq2$ with $K_X$ ample such that $GY(X)=0$; under these hypothesis it follows that:
\begin{enumerate}[a)]
\item $\fS$ is curve semistable;
\item $\fS_{\beta}$ is H-nef hence $\Omega^1_{\beta}$ is nef for every rational number $\displaystyle\beta\leq\frac{1}{n+1}$;
\item $\fS_{\beta}$ is H-ample hence $\Omega^1_{\beta}$ is ample for every rational number $\displaystyle\beta<\frac{1}{n+1}$;
\item for each irreducible projective curve $C$ on $X$ the inequality $\displaystyle2p_a(C)-2\geq\frac{1}{n+1}\int_CK_{X\vert C}$ holds;
\item $X$ does not contain irreducible projective curves of arithmetic genus neither $0$ nor $1$.
\end{enumerate}
These results can be inverted.
\begin{lemma}\label{lem4.1}
Let $X$ be a minimal smooth projective variety of general type of dimension $n\geq2$. If the Higgs bundle $\fS$ is semistable with respect to some polarization of $X$ and $\Delta(S)=0$ then:
\begin{enumerate}[a)]
\item $\fS$ is curve semistable;
\item\label{lem4.1.b} $\fS_{\beta}$ is H-nef hence $\Omega^1_{\beta}$ is nef for every rational number $\displaystyle\beta\leq\frac{1}{n+1}$;
\item\label{lem4.1.c} for each irreducible projective curve $C$ on $X$ the inequality $\displaystyle2p_a(C)-2\geq\frac{1}{n+1}\int_CK_{X\vert C}$ holds;
\item\label{lem4.1.d} $X$ does not contain irreducible projective curves of arithmetic genus neither $0$ nor $1$;
\item $K_X$ is ample;
\item\label{lem4.1.f} $\fS_{\beta}$ is H-ample hence $\Omega^1_{\beta}$ is ample for every rational number $\displaystyle\beta<\frac{1}{n+1}$.
\end{enumerate}
\end{lemma}
\begin{prf}
$\fS$ is curve semistable by Theorem \ref{th1.2}.\ref{th1.2.a} and applying Proposition \ref{prop2.2}.\ref{prop2.2.b}. Thus repeating the proof of Theorem \ref{th4.1}.\ref{th4.1.a}, one has the statement \ref{lem4.1.b}. Repeating the proof of Corollary \ref{cor4.1}, one has the statement \ref{lem4.1.c}. By the nefness of $K_X$, one has that $X$ does not contain irreducible projective curves of arithmetic genus $0$, hence $K_X$ is ample by \cite[Exercise 7.13.8]{D:O:2}. Thus and by statement \ref{lem4.1.c} one completes the proof of statement \ref{lem4.1.d}. Repeating the proof of Theorem \ref{th4.1}.\ref{th4.1.b} and applying Proposition \ref{prop2.1}.\ref{prop2.1.b}, one has the statement \ref{lem4.1.f}.
\end{prf}

\section{About algebraic hyperbolicity of smooth projective varieties\\ whose canonical bundle is ample}

\begin{definition}[cfr. {\cite[Definition 6.3.23.(i)]{L:RK}}]\label{def5.1}
A smooth projective variety $Y$ is \emph{algebraically hyperbolic} if there exists an ample divisor $L$ on $Y$ and $\epsilon\in\R_{>0}$ such that the arithmetic genus $p_a(C)$ of any projective curve $C$ on $Y$ satisfies the inequality
\begin{displaymath}
2p_a(C)-2\geq\epsilon\int_CL_{\vert C}.
\end{displaymath}
\end{definition}
\noindent By Corollary \ref{cor4.1}, without changing the notation of the previous definition, the following theorem holds.
\begin{theorem}\label{th3.5}
$X$ is algebraically hyperbolic choosing $L=K_X$ and $\displaystyle\epsilon=\frac{1}{n+1}$.
\end{theorem}
\begin{remark}
The previous theorem follows by Corollary \ref{cor3.3} and \cite[Theorem 6.3.26]{L:RK}. But in this way one does not obtain any estimate of $\epsilon$.
\end{remark}
\noindent Finally, by \cite[Proposition 9.8]{S:CT} one has another proof of the following theorem.
\begin{theorem}[{cfr. \cite[Theorem 4 and Remark (iii)]{Y:S-T}}]\label{th5.1}
If $\K=\C$ then $X$ is uniformized by $\B^n$ (the unit ball of $\C^n$).
\end{theorem}
\begin{remark}
By \cite[Example 2.1.2.2]{Z:L}, if $\K=\C$ then Corollary \ref{cor3.3} can be viewed as a consequence of the previous Theorem.
\end{remark}
\bigskip

\noindent{\bf Statement about competing or financial interests.} The author has no competing or financial interests to declare that are relevant to the content of this article.

\end{document}